# Improving Diffusion-based Data Augmentation with Inversion Spherical Interpolation

### v1

@@ 4.4 Ablation Study
Effectiveness of Each Component. Our Diff-II has two key components: Inversion Interpolation (II) and Two-stage Denoising (TD). We investigated the synthetic set of 5-shot Aircraft (same setting as Sec. 3.1 with ResNet) and reported: CLIP Score (Hessel et al., 2021) of the synthetic set; average LPIPS (Zhang et al., 2018) between images of the synthetic set, and classification accuracy. The CLIP score can reflect the faithfulness of the synthetic set while the LPIPS can indicate the diversity. As shown in Table 4, the first row (w/o both II and TD) directly denoise a random noise with a prompt without suffix in one stage(c.f., Sec. 3.3). We can see that: independently adding II or adding TD both can increase the LPIPS while nearly maintaining the CLIP Score. After adding both components together, the LPIPS further increased. This indicates that each component can significantly benefit the diversity with negligible harm to faithfulness, thus boosting higher accuracies.
Method land,land water,water land,water water,land Avg
@@
Real-Filter 38.87 34.72 38.25 31.64 35.87$(+1.17)$
Real-Guidance 39.49 33.64 39.80 29.78 35.68$(+0.98)$
Real-Mix 30.29 29.44 30.02 26.95 29.18$(-5.52)$
Da-Fusion 44.50 39.04 44.46 33.49 40.37$(+5.67)$
Diff-AUG 49.02 40.59 48.71 38.58 44.23$(+9.53)$
Diff-Mix 33.84 30.53 34.72 28.19 31.82$(-2.88)$
@@
Table 3: OOD classification. Results are averaged on 3 trials.
II TD CLIP Score ()($\uparrow$)( ↑ ) LPIPS ()($\uparrow$)( ↑ ) Acc ()($\uparrow$)( ↑ )
[…]
Table 4: Component Ablation. “II” is Inversion Interpolation and “TD” is Two-stage Denosing. “Acc” is the increase relative to no DA.
Figure 6: Influence of split ratio $s$. Except for the split ratio, all other settings and hyperparameters are the same with 5- shot CUB classification with ResNet50.
Split Ratio. Recall that in the Two-stage Denoising (c.f., Sec. 3.3), we have a split ratio $s$ to divide the denoising into two stages. To explore how the split ratio influences the synthetic data, we ablated it in Figure 6. This figure gives the curves of the CLIP score of the synthetic set and average LPIPS between images of the synthetic set changing with $s$. We can see that, with the increasing $s$, the CLIP Score decreases at a relatively slow rate while the LPIPS has a relatively large increase. By adjusting the value of $s$, a trade-off between faithfulness and diversity can be made.
Qutative Results. In Figure 7, we give some visualizations of Da-Fusion (Trabucco et al., 2023) and our Diff-II. We can see that the samples generated by DA-fusion lack diversity. In contrast, our Diff-II can generate samples with new context while maintaining the category characteristics.
Figure 7: Visualization Comparison. (a) Synthetic images of Da-fusion regarding different translation strengths. (b) Synthetic images of our Diff-II regarding different interpolation strengths.
@@ 5 Conclusion
In this work, we analyze current diffusion-based DA methods from a unified perspective, finding that they can either only improve the faithfulness of synthetic samples or only improve their diversity. To take both faithfulness and diversity into account, we propose Diff-II, a simple yet effective diffusion-based DA method. our Diff-II show that it significantly improves both the faithfulness and diversity of the synthetic samples, further improving classification models in data-scarce sceneries. In the future, we are going to: 1) extend this work into more general perception tasks, such as object detection, segmentation, or even video-domain tasks. 2) explore more effective DA methods that can better handle situations with extremely few training images.
@@
Limitations. Our method is less effective when some categories only have one training image. In that case, the interpolation can not be implemented because the interpolation operation is between two samples. We can see that for the long-tail classification task on CUB-LT (c.f., Table 2): as the imbalance factor gets larger (from 10 to 100), the gain of our method (compared with the second best one Diff-Mix)is getting smaller and smaller (from $5.8\%$ to $0.86\%$). This is because a higher imbalance factor means there are more categories that only have one training image.
Appendix A Appendix
This appendix is organized as follows:
  • Section A.1 gives the basic theory of spherical interpolation and the derived spherical extrapolation. Then we provide proof that the spherical interpolation result of two DDIM inversions is approximately standard normal distribution.
  • Section A.2 provides details of the comparison methods (c.f., Section 4), including Real-Filter (He et al., 2022), Real-Guidance (He et al., 2022), Da-Fusion (Trabucco et al., 2023), Real-Mix (Wang et al., 2024), Diff-AUG (Wang et al., 2024), Diff-Mix (Wang et al., 2024), CMO (Park et al., 2022) and CMO+DRW (Cao et al., 2019).
  • Section A.3 gives the implementation details of our Diff-II and the reproduction details of the comparison methods.
  • Section A.4 explain why interpolation results are better than random noise (c.f., Table 4).
  • Section A.5 shows additional results of experiments. First, we show the suffixes and predefined metaclasses of each dataset in Section A.5.1 and Section A.5.2 respectively. Then, we give the more ablation results of the split ratio (c.f., Section 3.3) in Section A.5.3. Finally, we give additional visualizations in Section A.5.4.
A.1 Theory of Spherical Interpolation
Spherical interpolation is a method used to interpolate between two points on a sphere or in a spherical space. The main idea behind spherical interpolation is to find a point along the shortest path on the sphere’s surface between two given points. The theory of spherical interpolation is grounded in spherical geometry and quaternion algebra as follows:
Shortest Path on Sphere: The shortest path between two points on the surface of a sphere is along the great circle that passes through both points. A great circle is any circle that divides the sphere into two equal hemispheres, like the equator or the meridians on a globe.
Interpolation Formula: Given two points on a sphere, represented by unit vectors $A$ and $B$, and an interpolation parameter $t\in[0,1]$, spherical interpolation calculates a third point $P$ along the great circle from $A$ to $B$ using the formula:
$P=\frac{\sin((1-t)\theta)A+\sin(t\theta)B}{\sin(\theta)},$ (9)
where $\theta$ is the angle between $A$ and $B$, found using the dot product $\cos(\theta)=A\cdot B$.
Quaternion Interpolation: When dealing with rotations in computer graphics, spherical interpolation can be applied using quaternions. Quaternions provide a way to represent orientations and rotations in three dimensions without the singularity and ambiguity problems of Euler angles. The interpolation of two quaternions $q_{1}$ and $q_{2}$ is given by:
$q=\frac{\sin((1-t)\theta)q_{1}+\sin(t\theta)q_{2}}{\sin(\theta)},$ (10)
where $\theta$ is the angle between the quaternions, computed as $\cos(\theta)=\text{Re}(q_{1}^{*}q_{2})$ (with $q_{1}^{*}$ being the conjugate of $q_{1}$). Spherical interpolation can smoothly interpolate rotations and directions, ensuring that the interpolated values remain on the sphere, and thus maintaining the integrity of the rotations or directional data.
Based on the above, we can easily derive the spherical interpolation $Z$ between two vectors ($I_{a}$ and $I_{b}$) of the same length:
$Z=\frac{sin((1-\lambda)\alpha)}{sin(\alpha)}I_{a}+\frac{sin(\alpha\lambda)}{sin(\alpha)}I_{b},\qquad\lambda\in[0,1],$ (11)
where $\lambda$ is the interpolation strength, $\alpha=arccos(\frac{I_{a}^{T}I_{b}}{(||I_{a}||||I_{b}||))}$ and $Z$ is the final interpolation result. Then we generalize to spherical extrapolation:
$Z=\frac{sin((1+\lambda)\alpha)}{sin(\alpha)}I_{a}-\frac{sin(\alpha\lambda)}{sin(\alpha)}I_{b},\qquad\lambda\in[0,2\pi/\alpha-1].$ (12)
Spherical extrapolation can expand the trajectory along the interpolation trajectory, increasing the interpolation range while still maintaining the integrity. Based on the periodicity of trigonometric functions, we can merge spherical interpolation and extrapolation:
$Z=\frac{sin((1+\lambda)\alpha)}{sin(\alpha)}I_{a}-\frac{sin(\alpha\lambda)}{sin(\alpha)}I_{b},\qquad\lambda\in[0,2\pi/\alpha].$ (13)
Then we give the proof that the spherical interpolation of two DDIM inversions is approximately standard normal distribution.
First, we consider that $I_{a}$ and $I_{b}$ are two DDIM inversions, which are approximately in standard normal distribution:
$I_{a}\sim N(\mu_{a},\sigma_{a}^{2}),\qquad\mu_{a}\simeq 0,\sigma_{a}\simeq 1.$ (14)
$I_{b}\sim N(\mu_{b},\sigma_{b}^{2}),\qquad\mu_{b}\simeq 0,\sigma_{b}\simeq 1.$ (15)
According to the superposition of normal distribution, we can get the distribution of $Z$:
$Z\sim N(\frac{sin((1+\lambda)\alpha)}{sin(\alpha)}\mu_{a}-\frac{sin(\alpha\lambda)}{sin(\alpha)}\mu_{b},(\frac{sin((1+\lambda)\alpha)}{sin(\alpha)})^{2}\sigma_{a}^{2}+(\frac{sin(\alpha\lambda)}{sin(\alpha)})^{2}\sigma_{b}^{2}).$ (16)
For the mean term of Eq. (16):
$\frac{sin((1+\lambda)\alpha)}{sin(\alpha)}\mu_{a}-\frac{sin(\alpha\lambda)}{sin(\alpha)}\mu_{b}\simeq 0.$ (17)
For the variance term, the $\alpha\simeq\pi/2$ due to $I_{a}$ and $I_{b}$ are two independent high-dimension vectors. Thus, $sin\alpha\simeq 1$ and $cos\alpha\simeq 0$. Then, we can simplify the variance term:
$(\frac{sin((1+\lambda)\alpha)}{sin(\alpha)})^{2}\sigma_{a}^{2}+(\frac{sin(\alpha\lambda)}{sin(\alpha)})^{2}\sigma_{b}^{2}$ (18)
$\simeq$ ${sin}^{2}((1+\lambda)\alpha)+{sin}^{2}(\alpha\lambda)$
$=$ ${sin}^{2}(\alpha+\alpha\lambda)+{sin}^{2}(\alpha\lambda)$
$=$ $(sin(\alpha)cos(\alpha\lambda)+cos(\alpha)sin(\alpha\lambda))^{2}+(sin(\alpha\lambda))^{2}$
$\simeq$ ${cos}^{2}(\alpha\lambda)+{sin}^{2}(\alpha\lambda)=1.$
Thus,
$Z\sim N(\mu,\sigma^{2}),\qquad\mu\simeq 0,\sigma\simeq 1.$ (19)
Proof completed.
A.2 Comparison methods
In this section, we introduce all the comparison methods of experiments.
Diffusion-based data augmentation methods:
  • Real-Filter (He et al., 2022): Directly generate some synthetic images with prompts containing their corresponding category labels. Then, leverage a pre-trained perception network to extract the features of both images of the original training set and synthetic images. Finally, filter all the synthetic images that are far from images of the original training set and only maintain those that are closed to the original training images.
  • Real-Guidance (He et al., 2022): Given an image from the original training set, add $T$ timesteps noise to the image and use the noised one to replace the random noise at the beginning of the generation. Finally, denoise it with a prompt containing its category label.
  • Da-Fusion (Trabucco et al., 2023): Firstly, set a few learnable token embeddings to learn an accurate concept for each category with the original training set. Then for a given image of the original training set, add random timesteps noise and denoise the noised image with a prompt containing its learned category concept.
  • Real-Mix (Wang et al., 2024): Given an image from the original training set, add random timesteps noise to the image. Then denoise the noised image with a prompt containing other-category labels. This will lead to a synthetic image with intermediate semantics between the two categories. Design a calculation mechanism to decide the soft label for this synthetic image.
  • Diff-AUG (Wang et al., 2024): Firstly, set a few learnable token embeddings and insert some learnable low-rank matrixes into the U-Net to learn an accurate concept for each category with the original training set. Then for a given image, add $T$ timesteps noise and denoise with a prompt containing its learned category concept.
  • Diff-Mix (Wang et al., 2024): Firstly, set a few learnable token embeddings and insert some learnable low-rank matrixes into the U-Net to learn an accurate concept for each category with the original training set. Then for a given image, add random timesteps noise and denoise with a prompt containing learned other-category concepts. This will lead to a synthetic image with intermediate semantics between the two categories. Design a calculation mechanism to decide the soft label for this synthetic image.
Long-tail classification methods:
  • CMO (Park et al., 2022): To balance the number of different categories’s training samples. CMO crops the objects from the rare-category images and pastes them to rich-category images to get some new images with rare-category objects and rich-category images’ backgrounds. These new images will be used to expand the rare-category images.
  • CMO+DRW (Cao et al., 2019): Except on oversampling-based CMO, DRW gives different weights to the loss of different categories. Specifically, the rare categories get a large loss weight while rich categories get a smaller loss weight.
A.3 Implementation details
In this section, we give all the implementation details of our Diff-II and reproduction details of comparison methods.
Details of our Diff-II:
  • Category concept learning: We follow the implementations of (Wang et al., 2024) [^3] https://github.com/Zhicaiwww/Diff-Mix.
  • Inversion interpolation: We use DDIM inversion (Song et al., 2020) with 25 steps and 1.0 guidance scale (Ho & Salimans, 2022) to calculate the inversion for each image. Then for each category, we randomly sample inversion pairs until the number of inversion pairs reaches five times the number of samples in the original training set. After that, we conduct spherical interpolation on these pairs with random strength $\lambda\in[0,1]$ (c.f., Section 3.2.2).
  • Two-stage denosing: We used BLIP-caption (Li et al., 2022) to get captions of all images. Then, we used GPT-4-turbo (Achiam et al., 2023) to summarize the captions into suffixes with the prompt:
    “I have a set of image captions that I want to summarize into objective descriptions that describe the scenes, actions, camera pose, zoom, and other image qualities present.
    My captions are: {captions}
    I want the output to be a $<=10$ of captions that describe a unique setting, of the form {prefix}.
    Here are 3 examples of what I want the output to look like:
    - {prefix} standing on a branch.
    - {prefix} flying in the sky with the Austin skyline in the background.
    - {prefix} playing in a river at night.
    Based on the above captions, the output should be:”
    Then, for each denoising, we randomly sampled a suffix for the first stage. For 5-shot classification, the split ratio was 0.3; for the 10-shot classification, the split ratio was 0.1; for the long-tail classification, the split ratio was 1.0. For the sample, we used the DDIM sampler with 25 steps and 7.5 guidance scale.
Details of comparison methods: For few-shot classification, we followed the reproduction implementations (i.e., the timesteps of adding noise) of (Wang et al., 2024). The translation strengths of Real-Guidance, Real-Mix, Da-Fusion, Diff-AUG, and Diff-Mix are 0.1, random one of [0.5, 0.7, 0.9], random one of [0.25, 0.5, 0.75, 1.0], 1.0, and random one of [0.5, 0.7, 0.9]. For long-tail classification, we directly report the results from (Wang et al., 2024).
Details of classifier training: For fairness, we used 0.5 as the replacement probability for all methods. Besides, we followed (Wang et al., 2024) for other settings and hyperparameters.
A.4 why Interpolation results better than random noise
Here we provide some explanations why starting with an interpolation result (the second row in Table 4) is better than a random noise (the first row in Table 4): Interpolation can not only sample some points in latent space that are not easy to sampled by standard normal distribution, but also the relative distance between these points will not be too close. This ensures the improvement of diversity. Besides, according to the characteristics of spherical interpolation, these points are in the position with relatively dense semantics of the pre-trained diffusion model, thus ensuring faithfulness. Therefore, the inversion interpolation results tend to generate more diverse samples than random Gaussian noise and can finally bootstrap better classification results.
A.5 Additional results
A.5.1 Suffixes
The suffixes of each dataset are listed as follows:
5-shot CUB:
  • standing on a tree branch.
  • flying around flowers.
  • standing on a post by the water.
  • flying over water.
  • standing on the ground.
  • swimming in the water.
  • sitting on a rock with a blue sky.
  • perched on a branch in a tree.
  • flying over water with wings spread.
  • perched on a tree branch.
10-shot CUB:
  • flying over water.
  • standing on the ground.
  • sitting on a rock.
  • swimming in water.
  • sitting on a bird feeder.
  • standing on the beach near water.
  • perched on a wire with a blue sky in the background.
  • standing on a branch with tall grass in the background.
  • flying in the sky with its wings spread.
5-shot Car:
  • parked on a city street.
  • on a white background.
  • parked in a lot with green trees in the background.
  • parked on a gravel road with mountains in the background.
  • driving down a tree-lined road.
  • parked on a black floor.
  • with its doors open.
  • charging at a station.
  • driving on a racing track.
10-shot Car:
  • parked on a road.
  • parked on a gravel road.
  • parked with trees in the background.
  • driving down a forested road.
  • driving on a dirt road in desert area.
  • on display at a show.
  • driving on a city street.
  • parked in a garage.
  • parked in front of a store with other cars.
5-shot Aircraft:
  • parked on the runway.
  • flying in the sky with the landing gear down.
  • landing with another plane in the background.
  • on the runway at an airport.
  • on the tarmac with mountains in the background.
  • flying in the air with the landing gear down.
  • parked in a hangar with the door open.
  • flying in the sky with palm trees in the background.
  • flying in the sky against a blue background.
  • lined up on the runway at the airport.
10-shot Aircraft:
  • flying in the sky with landing gear down.
  • taking off from the airport with a city in the background.
  • at the tarmac of an airport with a building in the background.
  • with passengers, flying in the sky.
  • propeller plane on a runway, with a Honeywell sign in the background.
  • airplane where workers are seen working on it in a hangar.
  • plane with a green stripe on the runway.
  • with people on board, floating in water.
  • with a red cross on its tail and landing gear.
  • jet on the runway with smoke coming out of it.
5-shot Pet:
  • lying on a pillow on the floor.
  • playing with a toy on the floor.
  • in the grass, looking at the camera with a leash.
  • on a window sill looking out.
  • sitting on a couch with a stuffed animal.
  • on a rock beside a person.
  • running in the grass with a frisbee.
10-shot Pet:
  • laying down in the grass.
  • sitting on brown leather furniture.
  • sitting on a couch with a dark background.
  • playing with a ball in the grass.
  • sitting on a windowsill, looking outdoors.
  • standing on a wooden deck.
  • sitting inside a cage.
  • sitting on a chair with its mouth open.
  • laying on a couch with a white background.
CUB-LT/IF=10:
  • flying over the ocean.
  • sitting on a rock by the water.
  • standing in the grass.
  • flying in the sky with its wings spread.
  • swimming in the water.
  • standing on a sandy beach.
  • sitting on a wire fence.
  • perched on a bird feeder in the snow.
  • standing on a tree stump.
CUB-LT/IF=20:
  • sitting on a rock in the water.
  • perched on a branch in a tree.
  • standing on the ground in the grass.
  • sitting on a post by the water.
  • standing on a ledge near the water.
  • flying in the sky with its wings spread.
  • sitting on a branch with a blurred background.
CUB-LT/IF=100:
  • flying in the sky with its wings spread.
  • standing on the ground in the dirt.
  • sitting on a branch of a tree.
  • swimming in the water.
  • perched on a hand in a grassy field.
  • standing on the shore of the water.
  • sitting on a branch of a tree.
  • sitting on a ledge by water.
  • standing in the water with its reflection.
Flower-LT/IF=10:
  • with water droplets on it.
  • growing in a garden.
  • in a close-up view.
  • with a bee on it.
  • in front of a water body.
  • against a brick wall.
  • with a butterfly on it.
  • with mixed colors in a bush.
Flower-LT/IF=20:
  • close-up with a dark background.
  • blooming in a garden.
  • growing in a pot.
  • floating on water in a pond.
  • arranged on a table.
  • with a bee on it in the garden.
  • on a tree.
  • in a field with a rocky surface.
  • against a blue sky.
  • with a blurry background in a field.
Flower-LT/IF=100:
  • with water droplets on it.
  • growing in a garden.
  • in a close-up view.
  • with a bee on it.
  • in front of a water body.
  • against a brick wall.
  • with a butterfly on it.
  • with mixed colors in a bush.
A.5.2 Prefdefined metaclasses
[…]
A.5.3 More ablation results for split ratio
Figure 8: Classification accuracy for different split ratios. Except for the split ratio, all other settings and hyperparameters are the same with 5-shot CUB classification with ResNet50.
We ablate the split ratio $s\in[0,1]$ in figure 8. we can see that: the value of $s$ will influence final classification accuracy. We get the best balance (when $s=0.3$) between faithfulness and diversity.
A.5.4 Additional Visualizations
Visualizations across different interpolation strength $\lambda$: As shown in Figure 9, we give our synthetic images regarding different interpolation strengths (c.f., Sec. 3.2.2). We can see that our Diff-II can generate samples with new context while maintaining the category concept characteristics. The interpolation strengths $\lambda$ can control the relative similarity between the synthetic sample and two samples of interpolation pair.
Synthetic images in few-shot classification and long-tail classification: we gave more synthetic images of our Diff-II used in few-shot and long-tail classification (c.f., Figure 10).
Figure 9: Synthetic images regarding different interpolation strengths
Figure 10: More synthetic images of our Diff-II in few-shot and long-tail classification.

### v2

@@ 4.4 Ablation Study
Figure 6: Visualization Comparison. (a) Synthetic images of Da-fusion regarding different translation strengths. (b) Synthetic images of our Diff-II regarding different interpolation strengths (The unit is $2\pi/\alpha$). Experientially, the interpolation type is extrapolation when the strength is in $[0,0.75]$, else interpolation.
Effectiveness of Each Component. We investigated the effectiveness of each component on the 5-shot Aircraft (same setting as Sec. 3.1 with ResNet) and reported: average LPIPS [40] between images of the synthetic set (which can reflect the diversity), and classification accuracy. As shown in Table 4, the first row indicates denoising a random noise with a templated prompt (e.g., “a photo of a Samoyed pet”). Since it lacks CL, the accuracy is quite poor, which indicates that for fine-grained datasets, concept learning is a crucial component. By comparing other rows, we can conclude that each component of our method (i.e., CL, SI, SE, TD) can improve the diversity and final classification performance. After combining all components, the LPIPS further increased, thus boosting higher accuracy.
Method L,L W,W L,W W,L Avg
@@
Real-Filter [12] 38.87 34.72 38.25 31.64 35.87$(+1.17)$
Real-Guidance [12] 39.49 33.64 39.80 29.78 35.68$(+0.98)$
Real-Mix [36] 30.29 29.44 30.02 26.95 29.18$(-5.52)$
Da-Fusion [33] 44.50 39.04 44.46 33.49 40.37$(+5.67)$
Diff-AUG [36] 49.02 40.59 48.71 38.58 44.23$(+9.53)$
Diff-Mix [36] 33.84 30.53 34.72 28.19 31.82$(-2.88)$
@@
Table 3: OOD classification. “L”, “W” represent “land” and “water”, respectively. Results are averaged on three trials.
CL LI SI SE TD LPIPS()($\uparrow$)( ↑ ) Acc()($\uparrow$)( ↑ )
49.9% 31.53
47.9% 38.74
27.5% 37.32
30.4% 38.37
35.4% 37.70
51.5% 40.49
50.0% 40.78
52.7% 42.15
Table 4: Components Ablation.CL” is Concept Learning, “LI” is Linear Interpolation, “SI” is Spherical Interpolation, “SE” is Spherical Extrapolation and “TD” is Two-stage Denosing.
Split Ratio. Recall that in the two-stage denoising (cf., Sec. 3.3), we have a split ratio $s$ to divide the denoising into two stages. To explore how the split ratio influences the synthetic data, we ablated it in Figure 7. This figure gives the curves of the CLIP score of the synthetic set and average LPIPS between images of the synthetic set changing with $s$. We can see that, with the increasing $s$, the CLIP score decreases at a relatively slow rate while the LPIPS has a relatively large increase. By adjusting $s$, a trade-off between faithfulness and diversity can be made.
Figure 7: Influence of split ratio $s$. Except for the split ratio, all other settings and hyperparameters are the same with 5- shot CUB classification with ResNet-50.
Qualitative Results. In Figure 6, we give some visualizations of Da-Fusion [33] and ours. We can see that the samples generated by DA-fusion lack diversity. In contrast, ours can generate samples with new context while maintaining the category characteristics.
@@ 5 Conclusion
In this work, we analyze current diffusion-based DA methods from a unified perspective, finding that they can either only improve the faithfulness of synthetic samples or only improve their diversity. To take both faithfulness and diversity into account, we propose Diff-II, a simple yet effective diffusion-based DA method. Our Diff-II shows that it significantly improves both the faithfulness and diversity of the synthetic samples, further improving classification models in data-scarce sceneries. In the future, we are going to: 1) Extend this work into more general perception tasks, such as object detection, segmentation, or even video-domain tasks. 2) Explore how to remove the dependency of captioner and only leverage LLMs to diversify the prompts used in two-stage denoising, making the augmentation more robust and can better handle out-of-distribution test sets.
@@
Limitations. Our method is less effective when some categories only have one training image. In that case, the interpolation can not be implemented because the interpolation operation is between two samples. We can see that for the long-tail classification task on CUB-LT (cf., Table 2): as the imbalance factor gets larger (from 10 to 100), the gain of our method (compared with the second best one Diff-Mix) is getting smaller and smaller (from $5.8\%$ to $0.86\%$). This is because a higher imbalance factor means there are more categories that only have one training image.